### v2

# Sums of four fourth power of primes

Abstract.
For any sufficiently large $\ell$, suppose that $\ell$ can be expressed as
$\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+p_{4}^{4},$
where $p_{1},p_{2},p_{3},p_{4}$ are primes.For such $\ell$, in this paper we will use circle method and sieves to prove that the proportion of $\ell$ in positive integers is at least $\frac{1}{27241.64}$ .
Key words and phrases:
Waring–Goldbach problem, Circle method, Sieve methods
1. Introduction
For any positive integers $m$ and $k$, the Warning-Goldbach Problem is to discuss every positive number $\ell_{m,k}$ can be expressed as the sum of the mthm-th powers of $k$ primes $p_{1},p_{2},\cdots,p_{k}$, i.e.
(1.1) $\ell_{m,k}=p_{1}^{m}+p_{2}^{m}+\cdots p_{k-1}^{m}+p_{k}^{m}.$
When $m=1,k=2$, this is the famous Goldbach’s Conjecture, the difficult problem, which has not been successfully solved so far. But when $m=1$, $k=3$, the case of three prime numbers has been proven, which is the famous theorem of three prime numbers[9]. Furthermore, for higher-order terms like $m=3,k=4$, some scholars have already obtained some results. Let $\mathcal{L}_{3,4}$ be the set of integers $\ell_{m,k}$ that can be written as the formula (1.1) when $m=3,k=4$ . In 1949, Roth[5] showed that
$\sum_{\ell_{3,4}\leqslant N\atop\ell_{3,4}\in\mathcal{L}_{3,4}}1\gg\frac{N}{\log^{8}N}.$
Futhermore, from 2001 to 2003, Ren[2][3]improved Roth’s theorem to the extent that
$\sum\limits_{\begin{subarray}{c}\ell_{3,4}\leqslant N\\ \ell_{3,4}\in\mathcal{L}_{3,4}\end{subarray}}1\geqslant\beta_{3,4}N,$
where $\beta_{3,4}=1/320$. In addition, Liu[4] improved this result and obtained $\beta_{3,4}=1/173.12$ in2012.
The authors are interested in this problem especially for more cases $m$ and $k$. In this paper we consider the case $m=4,k=4$ and try to say something about it. Let $\ell$ be a positive integer, suppose $\ell$ can be expressed as the sum of the four powers of four prime numbers, i.e.
(1.2) $\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+p_{4}^{4},$
define $\mathcal{L}$ as the set of all $\ell$, and assume that
(1.3) $\sum\limits_{\begin{subarray}{c}\ell_{\leqslant}N\\ \ell\in\mathcal{L}\end{subarray}}1\geqslant\beta N,$
Then we will consider the distribution of $\ell$. No one has yet conducted research on this problem, at least the authors have not see any relevant references. Therefore, we have the following theorem.
Theorem. $\beta=\frac{1}{27241.64}$ is acceptable in (1.3).
Note. Compared to the sums of powers of four cubes of primes, the density of sums of four fourth power of primes is much smaller, but if we expand the expression to the sums of more than four prime numbers, the density may become even higher, which is a problem worthy of further research.
2. Prerequisite Knowledge
Firstly, we introduce some definitions to prepare for the proof of the theorem.
Let NN be a large integer, $\delta$ be small enough and
$U=\left(\frac{N}{32(1+\delta)}\right)^{1/4},\quad V=U^{5/6}.$
Now, we define AA and $B$ as two sufficiently large integers, and $B$ is sufficiently large relative to AA. Define $\mathfrak{M}(q,a)$ as the interval $a/q-L^{B}/U^{4},a/q+L^{B}/U^{4}]$, then write $\mathfrak{M}$ for the union of all $\mathfrak{M}(q,a)$ with $1\leqslant a\leqslant q\leqslant L^{B}$ and $(a,q)=1.$ Obviously, $\mathfrak{M}(q,a)$ are disjoint. Define $\mathfrak{m}$ as the complement of $\mathfrak{M}$ in $L^{B}/U^{4},1+L^{B}/U^{4}]$.
For any real numbers $\lambda$ and $X$, define the following two integrals
$\Phi(\lambda,X)=\int\limits_{X}^{2X}e(u^{4}\lambda)\>du,\quad\Psi(\lambda,X)=\int\limits_{X}^{2X}\frac{e(u^{4}\lambda)}{\log u}\>du.$
Let
(2.1) $J(n)=\int_{-U^{4}L^{B}}^{U^{4}L^{B}}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\Big|\Psi(\lambda,U)\Big|^{2}\Big|\Psi(\lambda,V)\Big|^{4}e(-n\lambda)\>d\lambda.$
Define
(2.2) $\mathfrak{S}_{d}(n)=\sum_{q=1}^{\infty}T_{d}(n,q)\>,$
where
(2.3) $T_{d}(n,q)=\sum_{\overset{a=1}{(a,q)=1}}^{q}\frac{S(q,ad^{3})C^{3}(q,a)\overline{C^{4}(q,a)}}{q\varphi^{7}(q)}e\biggl(-\frac{an}{q}\biggr)$
and
(2.4) $S(q,a)=\sum_{m=1}^{q}e\biggl(\frac{am^{4}}{q}\biggr),\quad C(q,a)=\sum_{(m,q)=1}^{q}e\biggl(\frac{am^{4}}{q}\biggr).$
3. Some Lemmas
To prove the validity of the theorem, we also introduce the following lemmas:
Lemma 1. Let $0\leqslant|n|\leqslant N$, for each $m$ with $U<m\leqslant 2U$, denote by $R(m)$ the number of solutions of
$n=m^{4}+p_{2}^{4}+\cdots+p_{4}^{4}-p_{5}^{4}-\cdots-p_{8}^{4}$
with
$p_{2},p_{5},p_{6}\sim U,\quad p_{3},p_{4},p_{7},p_{8}\sim V.$
For $0<\xi<1/16$ and $D=N^{\xi}$, take $\mathfrak{S}_{d}(n)$ and $J(n)$ from the above (2.1), (2.2), define $E_{d}(n)$ as follows:
$\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0({\mathrm{mod}}~d)\end{subarray}}R(m)=\frac{\mathfrak{S}_{d}(n)}{d}J(n)+E_{d}(n).$
Then we have:
[…]
(ii) $J(n)$ is positive and satisfies
$J(n)\leqslant 11606V^{4}L^{-7}.$
(iii) For any complex numbers $\eta_{d}$ with $|\eta_{d}|\leqslant\tau(d)$, we hvae
$\sum_{d\leqslant D}\eta_{d}E_{d}(n)\ll V^{4}L^{-A}.$
Proof. Firstly, we prove (ii) of Lemma 1. From elementary estimation
$\int_{X^{4}}^{16X^{4}}e(\lambda u)\>du\leqslant\min(X^{4},|\lambda|^{-1})$
and integration by parts, we have
$\Phi(\lambda,X)=\frac{1}{4}\int_{X^{4}}^{16X^{4}}u^{-3/4}e(\lambda u)\,du\ll X^{-3}\min(X^{4},|\lambda|^{-1})$
and
$\Psi(\lambda,X)\ll X^{-3}\log^{-1}X\min(X^{4},|\lambda|^{-1}).$
We hope to calculate (2.1) by integrating the entire real axis, but this will result in the following error
$\ll\int_{U^{-4}L^{B}}^{\infty}|\Phi(\lambda,U)||\Psi(\lambda,U)|^{3}|\Psi(\lambda,V)|^{4}\>d\lambda$
$\ll U^{-12}V^{-12}L^{-7}\int_{U^{-4}L^{B}}^{\infty}\min(U^{4},|\lambda|^{-1})^{4}\min(V^{4},|\lambda|^{-1})^{4}\>d\lambda$
$\ll V^{4}L^{-3B}.$
By integral transformation, we have
$\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\big|\Psi(\lambda,U)\big|^{2}\big|\Psi(\lambda,V)\big|^{4}e(-n\lambda)\>d\lambda$
$=$ $\frac{1}{4}\int_{\mathcal{D}}\frac{d\nu_{1}\cdots d\nu_{4}du_{1}\cdots du_{3}}{\nu_{1}^{3/4}\cdots\nu_{4}^{3/4}u_{1}^{3/4}\cdots u_{4}^{3/4}\log\nu_{1}\cdots\log\nu_{4}\log u_{1}\cdots\log u_{3}},$
where
$\mathcal{D}=\left\{(v_{1},\ldots,v_{4},u_{1},\ldots,u_{3})\colon V^{4}\leqslant v_{1},\ldots,v_{4}\leqslant 16V^{4},\ U^{4}\leqslant u_{1},\ldots,u_{4}\leqslant 16U^{4}\right\}$
and $u_{4}=n+v_{1}+v_{2}-v_{3}-v_{4}+u_{1}+u_{2}-u_{3}.$
Then we get
(3.1) $\begin{split}J(n)=&\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}|\Psi(\lambda,U)|^{2}|\Psi(\lambda,V)|^{4}e(-n\lambda)d\lambda\\ &+\>O(V^{4}L^{-3B})\\ =&\frac{1}{4}\int_{V^{4}}^{16V^{4}}\>\frac{dv_{1}}{v_{1}^{3/4}\log v_{1}}\cdots\int_{V^{4}}^{16V^{4}}\>\frac{dv_{4}}{v_{4}^{3/4}\log v_{4}}\int_{U^{4}}^{16U^{4}}\>\frac{du_{1}}{u_{1}^{3/4}\log u_{1}}\\ &\times\int_{U^{4}}^{16U^{4}}\frac{du_{2}}{u_{2}^{3/4}\log u_{2}}\int_{\max(U^{4},x-16U^{4})}^{\min(16U^{4},x-U^{4})}\frac{du_{3}}{u_{3}^{3/4}(x-u_{3})^{3/4}\log u_{3}}\\ &+\>O(V^{4}L^{-3B}).\end{split}$
Let the last integral in the equation be denoted as II, where $x=n+v_{1}+v_{2}-v_{3}-v_{4}+u_{1}+u_{2}$, now we calculate the upper bound of II. Firstly, regarding $x\leqslant 2U^{4}$ or $x\geqslant 32U^{4}$, in both cases, the integral region does not exist, so we have $2U^{4}<x<32U^{4}$. In this case, we record as $u_{3}=xu$, then there is
$\begin{split}I&\leqslant(1+\varepsilon)x^{-1/2}L^{-1}\int_{\max(U^{4}/x,1-16U^{4}/x)}^{\min(16U^{4}/x,1-U^{4}/x)}u^{-3/4}(1-u)^{-3/4}\,du\\ &\leqslant\frac{1+\varepsilon}{\sqrt{2}}U^{-2}L^{-1}\int_{1/17}^{16/17}u^{-3/4}(1-u)^{-3/4}\,du\\ &\leqslant\frac{1+\varepsilon}{\sqrt{2}}U^{-2}L^{-1}I^{*},\end{split}$
where we record as $I^{*}=\int_{1/17}^{16/17}u^{-3/4}(1-u)^{-3/4}\,du$, then substituting it into (3.1) yields the following
$J(n)$ $\leq\frac{1+\varepsilon}{4\sqrt{2}}I^{*}U^{-2}L^{-1}\Bigg(\>\int_{U^{4}}^{16U^{4}}\>\frac{du}{u^{3/4}\log u}\Bigg)^{2}\Bigg(\>\int_{V^{4}}^{16V^{4}}\>\frac{dv}{v^{3/4}\log v}\Bigg)^{4}+O(V^{4}L^{-3B})$
$<\left(\frac{24}{5}\right)^{4}\frac{4}{\sqrt{2}}(1+\varepsilon)I^{*}V^{4}L^{-7}+O(V^{4}L^{-3B})$
$<11606V^{4}L^{-7},$
where we have used the estimate $I^{*}<7.73$. This proves (ii). For the proofs of (i) and (iii), we can refer to the methods of Brüdern [1] and Ren[2]. This proves Lemma 1.
Lemma 2. For $(d,6)=1$, we get
$\begin{split}\mathfrak{S}_{d}(n)=&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times\prod_{\begin{array}[]{c}p\nmid d\\ p\neq 2\end{array}}\{1+T_{1}(n,p)\}\prod_{p\mid d}\{1+T_{p}(n,p)\}.\end{split}$
Proof. According to the definition in (2.3), it can be obtained through a simple proof $\mathfrak{S}_{d}(n)$ which is an integral function, then there is
$\begin{split}\mathfrak{S}_{d}(n)&=\prod_{p}\{1+T_{p}(n,p)+T_{p}(n,p^{2})+\cdots\}\\ &=\prod_{p\nmid d}\{1+T_{1}(n,p)+T_{1}(n,p^{2})+\cdots\}\\ &\times\prod_{p\mid d}\{1+T_{p}(n,p)+T_{p}(n,p^{2})+\cdots\}.\end{split}$
For $p\nmid d$, we have
$\begin{split}T_{d}(n,p)&=\sum_{\overset{a=1}{(a,p)=1}}^{p}\frac{S(p,ad^{3})C^{3}(p,a)\overline{C^{4}(p,a)}}{p\varphi^{7}(p)}e\left(-\frac{an}{p}\right)\\ &=\sum_{\overset{a=1}{(a,p)=1}}^{p}\frac{S(p,a)C^{3}(p,a)\overline{C^{4}(p,a)}}{p\varphi^{7}(p)}e\left(-\frac{an}{p}\right)=T_{1}(n,p).\end{split}$
According to Lemma 4 in Hua[6], $C(p^{t},a)=0$($p\neq 2$ and $t\geq 2$ or $p=2,t\geq 5$), substituting back to the original equation yields the Lemma 2.
Lemma 3.Define $K(n,p)$ as the number of solutions to the following equation
$y_{2}^{4}+\cdots+y_{4}^{4}-y_{5}^{4}-\cdots-y_{8}^{4}\equiv n(\text{mod}~p)$
where $1\leq y_{i}\leq p-1$ $(1\leq i\leq 8)$, then we get
$pK(n,p)=(p-1)^{7}+E,$
where
$E=\begin{cases}\>-(p-1),&p\mid n,\\ \>1,&p\nmid n,\end{cases}~~~~~~~~~~~~~~\textit{if}~p\equiv 3({\mathrm{mod}}~4);$
and
$|E|\leqslant(3\sqrt{p}+1)^{5}(p-1)(3p+1),~~~\textit{if}~p\equiv 1({\mathrm{mod}}~4).$
It follows that $K(n,p)>0$ for $p\geq 13$.
Proof. From the definition of $K(n,p)$, we have
$pK(n,p)=\sum_{a=1}^{p}C^{3}(p,a)\overline{C^{4}(p,a)}e\left(-\frac{an}{p}\right)=(p-1)^{7}+E,$
where
$E=\sum_{a=1}^{p-1}C^{3}(p,a)\overline{C^{4}(p,a)}e\biggl(-\frac{an}{p}\biggr).$
When $p\equiv 3(\text{mod }4)$ and $(p,a)=1$, by Lemma 4.3 in Vaughan[7], we get $S(p,a)=0$ thus C(p,a)=1C(p,a)=-1, therefore
$E=-\sum_{a=1}^{p-1}e\left(-\frac{an}{p}\right)=\begin{cases}-(p-1),&p\mid n,\\ 1,&p\nmid n.\end{cases}$
and when $p\equiv 1(\text{mod }4)$, by Lemma 4.3 of Vaughan[7] again, we have $|C(p,a)|\leqslant 3\sqrt{p}+1$, and for
$\sum_{a=1}^{p-1}|C(p,a)|^{2}=\sum_{a=1}^{p}|C(p,a)|^{2}-(p-1)^{2}.$
Obiviously, $\sum_{a=1}^{p}|C(p,a)|^{2}$ can be expressed as $p$ times the number of solutions to equation $x^{4}\equiv y^{4}({\mathrm{mod~}}p)$, $1\leqslant x$, $y\leqslant p-1$. For $p\equiv 1({\mathrm{mod~}}4)$,
$\sum_{a=1}^{p}|C(p,a)|^{2}=4p(p-1),$
then
$\sum_{a=1}^{p-1}|C(p,a)|^{2}=(3p+1)(p-1),$
thus
$\sum_{a=1}^{p-1}|C^{3}(p,a)\overline{C^{4}(p,a)}|\leqslant(3\sqrt{p}+1)^{5}(p-1)(3p+1).$
A simple calculation shows that $K(n,p)>0$ for $p\geq 7$, $p\equiv 1(\text{mod }4)$and for all $p\equiv 3(\text{mod }4)$, therefore we have chosen $p\geq 13$. This proves Lemma 3.
From the similar method of Lemma 3, we can get Lemma 4.
Lemma 4. For $i=1,2$, let $H(n,p^{i})$ denote the number of solution of
$x^{3}+y_{2}^{3}+\cdots+y_{4}^{3}-y_{5}^{3}-\cdots-y_{8}^{3}\equiv n\pmod{p^{i}}$
where $1\leq x\leq p^{i},1\leq y_{j}<p^{i}$ and $(y_{j},p)=1$. Thus
$pH(n,p)=p(p-1)^{7}+E^{*},$
where
$E^{*}=0,~~~~~\textit{ if}~p\equiv 3(\text{mod }4);$
and
$|E^{*}|\leq 3\sqrt{p}(3\sqrt{p}+1)^{5}(p-1)(3p+1),~~\textit{if}~p\equiv 1(\text{mod }4).$
It follows that $H(n,p)>0$ for all pp.
Lemma 5. Definne the functions
$\phi_{0}(u)=\frac{2e^{\gamma_{0}}}{u}\mathrm{log}(u-1)\quad\text{and}\quad\phi_{1}(u)=\frac{2e^{\gamma_{0}}}{u},$
where $2\leq u\leq 3$. Suppose $\omega(d)$ is a multiplicative function of $d$ satisfying the conditions
$0\leqslant\omega(p)<p\quad\text{and}\quad\omega(p^{l})=1+O(p^{-1}),$
for each prime number $p$ and natural number $l$. Let $X$ be a real number with $X>3$, for $r(x)$ be a non-negative arithmetical function, we define
$E_{d}=\sum_{\begin{smallmatrix}P<x<2P\\ x\equiv 0\pmod{d}\end{smallmatrix}}r(x)-\frac{\omega(d)}{d}X.$
Let $U,V$ and $z$ be positive real parameters satisfying the inequality
$2\leqslant\frac{\log(UV)}{\log z}\leqslant 3.$
For any sequences $\{a_{m}\}$ and $\{b_{k}\}$ with
$|a_{m}|\leqslant 1\quad and\quad|b_{k}|\leqslant 1,$
one has
$\sum_{1\leq m\leq U}a_{m}\sum_{1\leq k\leq V}b_{k}E_{mk}\ll X(\log X)^{-2}.$
Then, we write
$W(z)=\prod_{p<z}\>(1-\omega(p)/p),$
one has the lower bound
$\sum_{\begin{subarray}{c}P\leq x<2P\\ (x,\Pi(z))=1\end{subarray}}r(x)>XW(z)\left(\phi_{0}\left(\frac{\log(UV)}{\log z}\right)+O\left((\log\log X)^{-1/50}\right)\right),$
and also the upper bound
$\sum_{\begin{subarray}{c}P\leq x<2P\\ (x,\Pi(z))=1\end{subarray}}r(x)<XW(z)\left(\phi_{1}\left(\frac{\log(UV)}{\log z}\right)+O\left((\log\log X)^{-1/50}\right)\right).$
Proof. See [8], Lemma 9.1.
Lemma 6. (Mertens’ theorem)For prime number $p$, $x\geq e$, positive integers $k$, $l$ with $(k,l)=1$ and $k\leq\ln^{A}{x}$ for any positive integer A, then
$\prod_{p\leq x,p\equiv l({\mathrm{mod~}k})}\left(1-\frac{1}{p}\right)=\frac{e^{-\frac{1}{\varphi(k)}\left[\gamma+\ln\frac{\varphi(k)}{k}\right]}}{\ln^{\frac{1}{\varphi(k)}}x}\left\{1+O\left(e^{-c\ln^{\frac{3}{5}}x}\right)\right\}$
where $c$ is a positive absolute constant, $\gamma$ is the Euler’s constant, and the constant in $O$ is independent of $x$.
Lemma 7. For $N/9<\ell\leqslant N$, define $r(\ell)$ as the number of $\ell$ can be expressed in the form of (1.2) with
$p_{1},p_{2}\sim U,\qquad p_{3},p_{4}\sim V\,.$
Then we have
$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)<bV^{4}L^{-8},$
where $b=3702014554$.
Proof. Denote $\mathcal{B}$ as the set of all prime numbers greater than 13, and define
$P(z)=\prod_{\begin{subarray}{c}p<z\\ p\in\mathcal{B}\end{subarray}}p.$
To prove Lemma 7, we hope to obtain an appropriate upper bound for the following equation
$\sum_{m\sim U\atop(m,P(z))=1}R(m).$
For this, we apply Lemma 5, there is no prime divisor beyond $\mathcal{B}$ for $d$, then we define
$\omega(d)=\mathfrak{S}_{d}(n)/\mathfrak{S}_{1}(n).$
Especially by Lemma 2, for $p\in\mathcal{B}$ we have
$\omega(p)=\frac{1+T_{p}(n,p)}{1+T_{1}(n,p)}.$
By (2.3), we get
$1+T_{p}(n,p)=\sum_{a=1}^{p}\frac{\overline{C(p,a)}|C(p,a)|^{6}}{(p-1)^{7}}e\Bigg(-\frac{an}{p}\Bigg)=p\frac{K(n,p)}{(p-1)^{7}}.$
hence
$\omega(p)=\frac{K(n,p)}{H(n,p)}.$
By Lemma 3 and Lemma 4, through simple calculations, for all $p\in\mathcal{B}$ and the positive integer $l$, we have
$0\leqslant\omega(p)<p,\quad\omega(p^{l})=1+O(p^{-1}).$
Now let $X=\mathfrak{S}_{1}(n)J(n)$, then by Lemma 1, we have
$\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0\mod d\end{subarray}}R(m)=\frac{\omega(d)}{d}X+E_{d}(n).$
Suppose $U_{0}\geqslant 1,\ V_{0}\geqslant 1,\ U_{0}V_{0}=D=N^{(1-\epsilon)/16}$ and $z=D^{1/2}$, then
$\frac{\log(U_{0}V_{0})}{\log z}=2.$
For any sequence of numbers $\{a_{m}\}$, $\{b_{k}\}$ with
$|a_{m}|\leq 1~~,~~|b_{k}|\leq 1,$
by Lemma 1, we have
$\sum_{1\leqslant m\leqslant U}a_{m}\sum_{1\leqslant k\leqslant V}b_{k}E_{mk}\ll\sum_{1\leqslant d\leqslant D}\tau(d)E_{d}\ll V^{4}L^{-A}.$
By Lemma 5, we get
$\sum\limits_{(m,P(z))=1}R(m)<e^{\gamma}(1+\varepsilon)J(n)\mathfrak{S}_{1}(n)W(z),$
where $\gamma$ denotes Euler’s constant, and
$W(z)=\prod_{\stackrel{{\scriptstyle p\in\mathcal{B}}}{{p<z}}}\biggl(1-\frac{\omega(p)}{p}\biggr).$
Actually, we have estimated the upper bound of $J(n)$ in Lemma 1. Next, we estimate the remaining part,
(3.2) $\begin{split}&\mathfrak{S}_{1}(n)W(z)\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))\\ &\times\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}(1+T_{1}(n,p))\biggl(1-\frac{K(n,p)}{H(n,p)}\biggr)\\ &\times\prod_{p\geqslant N^{(1-\varepsilon)/32}}(1+T_{1}(n,p))\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))\\ &\times\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1-\frac{1}{p}\right)\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\\ &\times\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}}{p(p-1)^{7}}\right).\end{split}$
For $n$ in the definition of $R(m)$, define $\rho(n)$ as the number of solutions to the following equation
$n=p_{1}^{4}+\cdots+p_{4}^{4}-p_{5}^{4}-\cdots-p_{8}^{4},$
where
$p_{1},p_{2},p_{5},p_{6}\sim U,\quad p_{3},p_{4},p_{7},p_{8}\sim V.$
We can naturally obtain
$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)\leqslant\rho(0)$
and
$\rho(n)\leq\sum_{m\sim U\atop(m,P(z))=1}R(m).$
Now we are calculating the various parts in (3.2). Since we are actually just trying to obtain a suitable upper bound for $\rho(0)$, thus we assume $n=0$ in the following text.
According to the definition (2.3), (2.4), by simple calculation we have $S(2,1)=0$, then
$T_{1}(n,2)=\sum_{\overset{a=1}{(a,2)=1}}^{2}\frac{S(2,a)C^{3}(2,a)\overline{C^{4}(2,a)}}{2\varphi^{7}(2)}e\biggl(-\frac{an}{2}\biggr)=0.$
For $q=4$, we have
$T_{1}(n,4)=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{S(4,a)C^{3}(4,a)\overline{C^{4}(4,a)}}{4\varphi^{7}(4)}e\biggl(-\frac{an}{4}\biggr),$
where
$S(4,a)=\sum_{m=1}^{4}e\biggl(\frac{am^{4}}{4}\biggr)=2+2e\biggl(\frac{a}{4}\biggl),C(4,a)=\sum_{(m,4)=1}^{4}e\biggl(\frac{am^{4}}{4}\biggr)=2e\biggl(\frac{a}{4}\biggl).$
Thus
$\begin{split}T_{1}(n,4)&=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{(2+2e(\frac{a}{4}))(2e(\frac{a}{4}))^{3}\overline{2e(\frac{a}{4})}^{4}}{4\varphi^{7}(4)}e\biggl(-\frac{a\times 0}{4}\biggr)\\ &=\frac{1}{2}\sum_{\overset{a=1}{(a,4)=1}}^{4}(1+e(-\frac{a}{4}))=\frac{1}{2}(2+2\cos{\frac{\pi}{2}})=1.\end{split}$
Similarly, we have
$T_{1}(n,8)=\frac{1}{2}(4+2\cos{\frac{\pi}{4}}+2\cos{\frac{3\pi}{4}})=2$
and
$T_{1}(n,16)=\frac{1}{2}(8+2\cos{\frac{\pi}{8}}+2\cos{\frac{3\pi}{8}}++2\cos{\frac{5\pi}{8}}+2\cos{\frac{7\pi}{8}})=4.$
Thus
$1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})=8.$
Also, for $q=3$, we have
$T_{1}(n,3)=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{S(3,a)C^{3}(3,a)\overline{C^{4}(3,a)}}{3\varphi^{7}(3)}e\biggl(-\frac{an}{3}\biggr),$
where
$S(3,a)=\sum_{m=1}^{3}e\biggl(\frac{am^{4}}{3}\biggr)=1+2e(\frac{a}{3}),C(4,a)=\sum_{(m,4)=1}^{4}e\biggl(\frac{am^{4}}{4}\biggr)=2e(\frac{a}{3}).$
then
$\begin{split}T_{1}(n,3)&=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{(1+2e(\frac{a}{3}))(2e(\frac{a}{3}))^{3}\overline{(2e(\frac{a}{3}))^{4}}}{3\varphi^{7}(3)}e\biggl(-\frac{a\times 0}{3}\biggr)\\ &=\frac{1}{3}(4+e(-\frac{1}{3})+e(-\frac{2}{3}))=\frac{1}{3}(4+2\cos{\frac{2\pi}{3}})=1.\end{split}$
Likewise, by a series of calculations we have $T_{1}(n,5)=\frac{1}{5}(16+2\cos{\frac{2\pi}{5}}+2\cos{\frac{4\pi}{5}})=\frac{1}{5}(16+\frac{\sqrt{5}-1}{2}-\frac{\sqrt{5}+1}{2})=3$, $T_{1}(n,7)=\frac{8}{729}$ and $T_{1}(n,11)=\frac{324}{171875}.$
Furthermore, for
$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right),$
by Lemma 3 and Lemma 4, for $p\equiv 3(\text{mod }4)$ we have
$|E-E^{*}|\leq p-1,$
then
$\begin{split}\prod_{\overset{13\leqslant p<N^{(1-\varepsilon)/32}}{p\equiv 3(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)&\leq\prod_{\overset{p\geqslant 13}{p\equiv 3(\mathrm{mod~4})}}\left(1+\frac{1}{\left(p-1\right)^{7}}\right)\\ &<1\>+\>\sum_{n\geqslant 12}\>\frac{1}{n^{7}}<1\>+\>12^{-6},\end{split}$
and for $p\equiv 1(\text{mod }4)$, we have
$|E-E^{*}|\leq(3\sqrt{p}+1)^{6}(p-1)(3p+1).$
then for $p\geq 101$, by $1+x\leq e^{x}$, we have
$\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}e^{\frac{E^{*}-E}{\left(p-1\right)^{8}}}=e^{\sum_{p\geq 101}{\frac{E^{*}-E}{\left(p-1\right)^{8}}}},$
where
$\sum_{p\geq 101}{\frac{|E^{*}-E|}{\left(p-1\right)^{8}}}\leq 0.305,$
thus
$\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq 1.357.$
In addition,
$\prod_{\overset{13\leq p\leq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq 11.773.$
Therefore we have
$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)<15.976$
Furthermore, by Lemma 6, we have the estimate
$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1-\frac{1}{p}\right)<308e^{-\gamma}(1+\varepsilon)L^{-1}.$
It is not difficult to find that
$\frac{|E^{*}|}{p(p-1)^{7}}<\frac{1}{p^{2}}$
for $p$ large enough. Hence
$\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}}{p(p-1)^{7}}\right)\leqslant\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{1}{p^{2}}\right)<1+\varepsilon.$
To sum up,
$\mathfrak{S}_{1}(n)W(z)<318974.199e^{-\gamma}L^{-1}$
then
$\sum_{m\sim U\atop(m,P(z))=1}R(m)<bV^{4}L^{-8},$
where $b=3702014554$.
Naturally, for $n=0$,
$\rho(0)\leq\sum_{m\sim U\atop(m,P(z))=1}R(m)$
and
$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)\leq\rho(0),$
therefore
$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)<bV^{4}L^{-8}.$
We have proved Lemma 5.
4. Proof of the Theorem
In this part, we will prove Theorem.
By the prime number theorem, for
$U=\left(\frac{N}{32(1+\delta)}\right)^{1/4},\quad V=U^{5/6},$
we have
(4.1) $\begin{split}\sum_{N/9<\ell\leqslant N}r(\ell)&\geqslant\sum_{p_{1}\sim U}1\sum_{p_{2}\sim U}1\sum_{p_{3}\sim V}1\sum_{p_{4}\sim V}1\geqslant(1-\varepsilon)\frac{U^{2}V^{2}}{\log^{2}U\log^{2}V}\\ &\geqslant\left(\frac{96}{5}\right)^{2}(1-\varepsilon)U^{2}V^{2}L^{-4}.\end{split}$
then by Cauchy’s inequality and Lemma 5, we have
(4.2) $\begin{split}\left\{\sum_{N/9<\ell\leqslant N}r(\ell)\right\}^{2}&\leqslant\left\{\sum_{\begin{subarray}{c}N/9<\ell\leqslant N\\ r(\ell)>0\end{subarray}}1\right\}\left\{\sum_{N/9<\ell\leqslant N}r^{2}(\ell)\right\}\\ &\leqslant bV^{4}L^{-8}\left\{\sum_{N/9<\ell\leqslant N}1\right\}.\end{split}$
From (4.1), (4.2), we have
$\sum_{N/9<\ell\leqslant N\atop r(\ell)>0}1>\frac{(1-\varepsilon)^{2}}{b}\biggl(\frac{96}{5}\biggr)^{4}U^{4}>\frac{1}{27241.64}N.$
The proof of our Theorem is now complete.
$\square$

### v5

@@
-Sums of four fourth power of primes
+Sums of eight fourth power of primes
 Abstract.
 For any sufficiently large $\ell$, suppose that $\ell$ can be expressed as
-$\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+p_{4}^{4},$
-where $p_{1},p_{2},p_{3},p_{4}$ are primes.For such $\ell$, in this paper we will use circle method and sieves to prove that the proportion of $\ell$ in positive integers is at least $\frac{1}{27241.64}$ .
+$\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+\cdots+p_{8}^{4},$
+where $p_{1},p_{2},p_{3},\cdots,p_{8}$ are primes.For such $\ell$, in this paper we will use circle method and sieves to prove that the proportion of $\ell$ in positive integers is at least $\frac{1}{414.465}$ .
 Key words and phrases:
@@
 1. Introduction
-For any positive integers $m$ and $k$, the Warning-Goldbach Problem is to discuss every positive number $\ell_{m,k}$ can be expressed as the sum of the mthm-th powers of $k$ primes $p_{1},p_{2},\cdots,p_{k}$, i.e.
+For any positive integers $m$ and $k$, the Waring-Goldbach Problem is to discuss every positive number $\ell_{m,k}$ can be expressed as the sum of the $m$-thth powers of $k$ primes $p_{1},p_{2},\cdots,p_{k}$, i.e.
 (1.1) $\ell_{m,k}=p_{1}^{m}+p_{2}^{m}+\cdots p_{k-1}^{m}+p_{k}^{m}.$
@@
 But when $m=1$, $k=3$, the case of three prime numbers has been proven, which is the famous theorem of three prime numbers[9].
-Furthermore, for higher-order terms like $m=3,k=4$, some scholars have already obtained some results.
+Furthermore, for higher-order terms $m\geq 1$, scholars hope to obtain the smallest $k$ that enables all positive integers to be expressed in the form of (1.1), and denote this smallest $k$ as $G(m)$.
+In 1770, Waring deduced from a finite set of evidence that every positive integer is the sum of four squares, nine cubes, nineteen fourth powers, and so on.
+In 1770, Lagrange proved the existence of $G(2)$, and in the following 139 years, proofs of existence were obtained for $m=3,4,5,6,7,8,10$.
+In 1909, Hilbert used induction to prove the existence of G(m)G(m)for mm[13].
+In addition, for those $m,k$ that cannot express all positive integers in the form of (1.1), scholars also hope to study the density of positive integers that can be expressed in the form of (1.1) among all positive integers in this case.
+For example $m=3,k=4$, some scholars have already obtained some results.
 Let $\mathcal{L}_{3,4}$ be the set of integers $\ell_{m,k}$ that can be written as the formula (1.1) when $m=3,k=4$ .
@@
 $\sum_{\ell_{3,4}\leqslant N\atop\ell_{3,4}\in\mathcal{L}_{3,4}}1\gg\frac{N}{\log^{8}N}.$
-Futhermore, from 2001 to 2003, Ren[2][3]improved Roth’s theorem to the extent that
+Furthermore, from 2001 to 2003, Ren[2][3]improved Roth’s theorem to the extent that
 $\sum\limits_{\begin{subarray}{c}\ell_{3,4}\leqslant N\\ \ell_{3,4}\in\mathcal{L}_{3,4}\end{subarray}}1\geqslant\beta_{3,4}N,$
@@
 The authors are interested in this problem especially for more cases $m$ and $k$.
-In this paper we consider the case $m=4,k=4$ and try to say something about it.
-Let $\ell$ be a positive integer, suppose $\ell$ can be expressed as the sum of the four powers of four prime numbers, i.e.
-(1.2) $\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+p_{4}^{4},$
+In this paper we consider the case $m=4,k=7$ and try to say something about it.
+Let $\ell$ be a positive integer, suppose $\ell$ can be expressed as the sum of the fourth powers of eight prime numbers, i.e.
+(1.2) $\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+\cdots+p_{8}^{4},$
 define $\mathcal{L}$ as the set of all $\ell$, and assume that
@@
 Therefore, we have the following theorem.
-Theorem. $\beta=\frac{1}{27241.64}$ is acceptable in (1.3).
+Theorem. $\beta=\frac{1}{414.465}$ is acceptable in (1.3).
 Note.
-Compared to the sums of powers of four cubes of primes, the density of sums of four fourth power of primes is much smaller, but if we expand the expression to the sums of more than four prime numbers, the density may become even higher, which is a problem worthy of further research.
+Compared to the sums of four cubes of primes, the density of sums of eight fourth power of primes is smaller.
+For cases with a relatively small number, the results obtained by the method in this paper are not ideal, but if we expand the expression to the sums of more than eight prime numbers, the density may become even higher, which is a problem worthy of further research.
 2. Prerequisite Knowledge
 Firstly, we introduce some definitions to prepare for the proof of the theorem.
-Let NN be a large integer, $\delta$ be small enough and
-$U=\left(\frac{N}{32(1+\delta)}\right)^{1/4},\quad V=U^{5/6}.$
+Let $N$ be a large integer, $\delta_{0}$ be small enough and
+$U=\left(\frac{N}{64(1+\delta_{0})}\right)^{1/4},\quad V=U^{7/8}.$
 Now, we define AA and $B$ as two sufficiently large integers, and $B$ is sufficiently large relative to AA.
@@
 Let
-(2.1) $J(n)=\int_{-U^{4}L^{B}}^{U^{4}L^{B}}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\Big|\Psi(\lambda,U)\Big|^{2}\Big|\Psi(\lambda,V)\Big|^{4}e(-n\lambda)\>d\lambda.$
+(2.1) $J(n)=\int_{-U^{-4}L^{B}}^{U^{-4}L^{B}}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\Big|\Psi(\lambda,U)\Big|^{6}\Big|\Psi(\lambda,V)\Big|^{8}e(-n\lambda)\>d\lambda.$
 Define
@@
 where
-(2.3) $T_{d}(n,q)=\sum_{\overset{a=1}{(a,q)=1}}^{q}\frac{S(q,ad^{3})C^{3}(q,a)\overline{C^{4}(q,a)}}{q\varphi^{7}(q)}e\biggl(-\frac{an}{q}\biggr)$
+(2.3) $T_{d}(n,q)=\sum_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\frac{S(q,ad^{4})C^{7}(q,a)\overline{C^{8}(q,a)}}{q\varphi^{15}(q)}e\biggl(-\frac{an}{q}\biggr)$
 and
-(2.4) $S(q,a)=\sum_{m=1}^{q}e\biggl(\frac{am^{4}}{q}\biggr),\quad C(q,a)=\sum_{(m,q)=1}^{q}e\biggl(\frac{am^{4}}{q}\biggr).$
+(2.4) $S(q,a)=\sum_{m=1}^{q}e\biggl(\frac{am^{4}}{q}\biggr),\quad C(q,a)=\sum_{\begin{subarray}{c}q=1\\ (q,m)=1\end{subarray}}^{q}e\biggl(\frac{am^{4}}{q}\biggr).$
 3. Some Lemmas
 To prove the validity of the theorem, we also introduce the following lemmas:
+Lemma 1. The number S of solutions of
+(3.1) $x_{1}^{4}+y_{1}^{4}+y_{2}^{4}=x_{2}^{4}+y_{3}^{4}+y_{4}^{4}$
+with $U<x\leq 2U,U^{7/8}<y\leq 2U^{7/8}$ satisfies
+$S\ll U^{25/8}.$
+Proof. For $x_{1}=x_{2}$, we can easily calculate that the number $S$ of solutions of (3.1) satisfies
+$S\ll U^{25/8}.$
+For $x_{1}\neq x_{2}$, base on symmetry, we assume $x_{1}<x_{2}$ and write $x_{2}=x_{1}+h$. Then (3.1) becomes
+$h(4x_{1}^{3}+6x_{1}^{2}h+4x_{1}h^{2}+h^{3})=y_{1}^{4}+y_{2}^{4}-y_{3}^{4}+y_{4}^{4}.$
+Since $y_{1}^{4}+y_{2}^{4}\leq 32U^{7/2}$ and $x_{1}^{3}>U^{3}$ it follows that $h<8U^{1/2}$.
+Let
+$F_{h}(\alpha)\>=\>\sum_{U<x\>\leqslant\>2U}e\bigl(\alpha h(4x_{1}^{3}+6x_{1}^{2}h+4x_{1}h^{2}+h^{3})\bigr)\>,$
+$G(\alpha)=\sum_{h<8U^{1/2}}F_{h}(\alpha)\>,$
+$f(\alpha)\>=\>\sum_{U^{7/8}<y\>\leqslant\>2U^{7/8}}e(\alpha y^{4})\>.$
+Then
+$S\leqslant\int_{0}^{1}G(\alpha)|f(\alpha)|^{4}d\alpha.$
+We hope to obtain an upper bound for $G(\alpha)$. Suppose that $|\alpha-a/q|<q^{-2}$ and $(a,q)=1$. By Cauchy’ s inequality
+$|G(\alpha)|^{2}\>\ll U^{1/2}\>\sum_{h\><\>8U^{1/2}}|F_{h}(\alpha)|^{2}.$
+Moreover, write $y=x+h_{1}$ and we have
+$|F_{h}(\alpha)|^{2}\>$ $=\sum_{U<x\>\leqslant\>2U}\sum_{U<y\>\leqslant\>2U}e\bigl(\alpha h(4x_{1}^{3}+6x_{1}^{2}h+4x_{1}h^{2}-4y_{1}^{3}-6y_{1}^{2}h-4y_{1}h^{2}\bigr)$
+$=\>\sum_{|h_{1}|<U}\sum_{\max(U,U-h_{1})<x\leqslant\>\min(2U,2U-h_{1})}e\bigl(\alpha h(12h_{1}x^{2}+12h_{1}(h+h_{1})x+4h_{1}^{3}+6hh_{1}^{2}+4h^{2}h_{1})\bigr).$
+Record the innermost sum formula as $T(h_{1})$. Use Cauchy’ s inequality again
+$|F_{h}(\alpha)|^{2}\>\leq U^{1/2}\biggl(\sum_{|h_{1}|<U}|T(h_{1})|^{2}\biggl)^{1/2}.$
+Similarly, write $y^{\prime}=x^{\prime}+h_{1}$, then
+$|T(h_{1})|^{2}$ $=\sum_{|h_{2}|<U}\sum_{\max(U,U-h_{2})<x^{\prime}\leqslant\>\min(2U,2U-h_{2})}e(\alpha hh_{1}h_{2}(24x+12(h+h_{1}+h_{2})))$
+[…]
+Hence
+$|F_{h}(\alpha)|^{2}$ $\ll U^{1/2}\biggl(\sum_{|h_{1}|<U}\bigg(U+\sum_{0<h_{2}<U}\min(U,||24\alpha hh_{1}h_{2}||^{-1})\bigg)\biggl)^{1/2}$
+[…]
+By Lemma 2.2 of Vaughan[7], we have
+(3.2) $|G(\alpha)|^{2}\ll U^{5/2+\epsilon}+U^{13/4+\epsilon}q^{-1/2}+U^{11/4+\epsilon}+U^{3/2+\epsilon}q^{1/2}$
+Let $\mathfrak{M}^{\prime}(q,a)$ denote the interval $a/q-q^{-1}U^{-2},a/q+q^{-1}U^{-2}]$ and $\mathcal{U}=(P^{-2},1+P^{-2}]$.
+We may suppose that $U\geqslant 4$.
+Then the $\mathfrak{M}^{\prime}(q,a)$ with $1\leqslant a\leqslant q\leqslant U$, $(a,q)=1$ are disjoint and contained in $\mathcal{U}$.
+Let $\mathfrak{M}$ be the union of the $\mathfrak{M}^{\prime}(q,a)$ with $1\leqslant a\leqslant q\leqslant U$, $(a,q)=1$, and let $m^{\prime}=\mathcal{U}\backslash\mathfrak{M}^{\prime}$.
+Then
+$S\leqslant\>\int_{{}_{\mathfrak{M}^{\prime}}}G(\alpha)|f(\alpha)|^{4}d\alpha\>+\>\int_{{}_{m^{\prime}}}G(\alpha)|f(\alpha)|^{4}\>d\alpha.$
+By (3.2), when $q>U$, there is $\alpha\notin\mathfrak{M}^{\prime}$ and
+$G(\alpha)\ll P^{11/8+\epsilon}.$
+Hence, by Lemma 2.5 of Vaughan[7],
+$\int_{{}_{m^{\prime}}}G(\alpha)|f(\alpha)|^{4}\>d\alpha\ll P^{25/8}.$
+For $\alpha\in\mathfrak{M}^{\prime}$, there is $1\leqslant a\leqslant q\leqslant U$, then by Lemma 6.3 of Vaughan[7]
+$G(\alpha)\ll U^{13/8}q^{-1/4},$
+$f(\alpha)\ll U^{7/8}q^{-1/4}(1+U^{7/2}|\alpha-a/q|)^{-1}.$
+Therefore
+$\int_{{}_{\mathfrak{M}^{\prime}}}G(\alpha)|f(\alpha)|^{4}d\alpha$ $\ll\sum_{q\leq U}\sum_{\stackrel{{\scriptstyle a=1}}{{(a,q)=1}}}^{q}U^{13/8+7/2}q^{-1-1/4}\int_{\mathfrak{M}^{\prime}(q,a)}\left(1+U^{7/2}\bigg|\alpha-\frac{a}{q}\bigg|\right)^{-4}d\alpha$
+$\ll U^{19/8}$
+This proves Lemma 1.
 Lemma 2. Let $0\leqslant|n|\leqslant N$, for each $m$ with $U<m\leqslant 2U$, denote by $R(m)$ the number of solutions of
-$n=m^{4}+p_{2}^{4}+\cdots+p_{4}^{4}-p_{5}^{4}-\cdots-p_{8}^{4}$
+$n=m^{4}+p_{2}^{4}+\cdots+p_{8}^{4}-p_{9}^{4}-\cdots-p_{16}^{4}$
 with
-$p_{2},p_{5},p_{6}\sim U,\quad p_{3},p_{4},p_{7},p_{8}\sim V.$
-For $0<\xi<1/16$ and $D=N^{\xi}$, take $\mathfrak{S}_{d}(n)$ and $J(n)$ from the above (2.1), (2.2), define $E_{d}(n)$ as follows:
+$p_{2},p_{3},p_{4},p_{9},p_{10},p_{11},p_{12}\sim U,\quad p_{5},p_{6},p_{7},p_{8},p_{13},p_{14},p_{15},p_{16}\sim V.$
+For $0<\xi<9/25$ and $D=N^{\xi}$, take $\mathfrak{S}_{d}(n)$ and $J(n)$ from the above (2.1), (2.2), define $E_{d}(n)$ as follows:
 $\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0({\mathrm{mod}}~d)\end{subarray}}R(m)=\frac{\mathfrak{S}_{d}(n)}{d}J(n)+E_{d}(n).$
@@
 (ii) $J(n)$ is positive and satisfies
-$J(n)\leqslant 11606V^{4}L^{-7}.$
+$J(n)\leqslant KU^{4}V^{8}L^{-15}.$
+where $K=4888799.222$
 (iii) For any complex numbers $\eta_{d}$ with $|\eta_{d}|\leqslant\tau(d)$, we hvae
-$\sum_{d\leqslant D}\eta_{d}E_{d}(n)\ll V^{4}L^{-A}.$
+$\sum_{d\leqslant D}\eta_{d}E_{d}(n)\ll U^{4}V^{8}L^{-A}.$
 Proof. Firstly, we prove (ii) of Lemma 2. From elementary estimation
@@
 We hope to calculate (2.1) by integrating the entire real axis, but this will result in the following error
-$\ll\int_{U^{-4}L^{B}}^{\infty}|\Phi(\lambda,U)||\Psi(\lambda,U)|^{3}|\Psi(\lambda,V)|^{4}\>d\lambda$
-$\ll U^{-12}V^{-12}L^{-7}\int_{U^{-4}L^{B}}^{\infty}\min(U^{4},|\lambda|^{-1})^{4}\min(V^{4},|\lambda|^{-1})^{4}\>d\lambda$
-$\ll V^{4}L^{-3B}.$
+$\ll\int_{U^{-4}L^{B}}^{\infty}|\Phi(\lambda,U)||\Psi(\lambda,U)|^{7}|\Psi(\lambda,V)|^{8}\>d\lambda$
+$\ll U^{-24}V^{-24}L^{-15}\int_{U^{-4}L^{B}}^{\infty}\min(U^{4},|\lambda|^{-1})^{6}\min(V^{4},|\lambda|^{-1})^{8}\>d\lambda$
+$\ll U^{4}V^{8}L^{-5B}.$
 By integral transformation, we have
-$\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\big|\Psi(\lambda,U)\big|^{2}\big|\Psi(\lambda,V)\big|^{4}e(-n\lambda)\>d\lambda$
-$=$ $\frac{1}{4}\int_{\mathcal{D}}\frac{d\nu_{1}\cdots d\nu_{4}du_{1}\cdots du_{3}}{\nu_{1}^{3/4}\cdots\nu_{4}^{3/4}u_{1}^{3/4}\cdots u_{4}^{3/4}\log\nu_{1}\cdots\log\nu_{4}\log u_{1}\cdots\log u_{3}},$
+$\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\big|\Psi(\lambda,U)\big|^{6}\big|\Psi(\lambda,V)\big|^{8}e(-n\lambda)\>d\lambda$
+$=$ $\frac{1}{4}\int_{\mathcal{D}}\frac{d\nu_{1}\cdots d\nu_{8}du_{1}\cdots du_{7}}{\nu_{1}^{3/4}\cdots\nu_{8}^{3/4}u_{1}^{3/4}\cdots u_{8}^{3/4}\log\nu_{1}\cdots\log\nu_{8}\log u_{1}\cdots\log u_{7}},$
 where
-$\mathcal{D}=\left\{(v_{1},\ldots,v_{4},u_{1},\ldots,u_{3})\colon V^{4}\leqslant v_{1},\ldots,v_{4}\leqslant 16V^{4},\ U^{4}\leqslant u_{1},\ldots,u_{4}\leqslant 16U^{4}\right\}$
-and $u_{4}=n+v_{1}+v_{2}-v_{3}-v_{4}+u_{1}+u_{2}-u_{3}.$
+$\mathcal{D}=\left\{(v_{1},\ldots,v_{8},u_{1},\ldots,u_{7})\colon V^{4}\leqslant v_{1},\ldots,v_{8}\leqslant 16V^{4},\ U^{4}\leqslant u_{1},\ldots,u_{7}\leqslant 16U^{4}\right\}$
+and $u_{8}=n+v_{1}+v_{2}+v_{3}+v_{4}-v_{5}-v_{6}-v_{7}-v_{8}+u_{1}+u_{2}+u_{3}+u_{4}-u_{5}-u_{6}-u_{7}.$
 Then we get
-(3.1) $\begin{split}J(n)=&\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}|\Psi(\lambda,U)|^{2}|\Psi(\lambda,V)|^{4}e(-n\lambda)d\lambda\\ &+\>O(V^{4}L^{-3B})\\ =&\frac{1}{4}\int_{V^{4}}^{16V^{4}}\>\frac{dv_{1}}{v_{1}^{3/4}\log v_{1}}\cdots\int_{V^{4}}^{16V^{4}}\>\frac{dv_{4}}{v_{4}^{3/4}\log v_{4}}\int_{U^{4}}^{16U^{4}}\>\frac{du_{1}}{u_{1}^{3/4}\log u_{1}}\\ &\times\int_{U^{4}}^{16U^{4}}\frac{du_{2}}{u_{2}^{3/4}\log u_{2}}\int_{\max(U^{4},x-16U^{4})}^{\min(16U^{4},x-U^{4})}\frac{du_{3}}{u_{3}^{3/4}(x-u_{3})^{3/4}\log u_{3}}\\ &+\>O(V^{4}L^{-3B}).\end{split}$
-Let the last integral in the equation be denoted as II, where $x=n+v_{1}+v_{2}-v_{3}-v_{4}+u_{1}+u_{2}$, now we calculate the upper bound of II.
+(3.3) $\begin{split}J(n)=&\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}|\Psi(\lambda,U)|^{6}|\Psi(\lambda,V)|^{8}e(-n\lambda)d\lambda\\ &+\>O(U^{4}V^{8}L^{-5B})\\ =&\frac{1}{4}\int_{V^{4}}^{16V^{4}}\>\frac{dv_{1}}{v_{1}^{3/4}\log v_{1}}\cdots\int_{V^{4}}^{16V^{4}}\>\frac{dv_{4}}{v_{8}^{3/4}\log v_{8}}\int_{U^{4}}^{16U^{4}}\>\frac{du_{1}}{u_{1}^{3/4}\log u_{1}}\\ &\times\int_{U^{4}}^{16U^{4}}\frac{du_{2}}{u_{2}^{3/4}\log u_{2}}\cdots\int_{\max(U^{4},x-16U^{4})}^{\min(16U^{4},x-U^{4})}\frac{du_{7}}{u_{7}^{3/4}(x-u_{7})^{3/4}\log u_{7}}\\ &+\>O(U^{4}V^{8}L^{-5B}).\end{split}$
+Let the last integral in the equation be denoted as $I$, where $x=n+v_{1}+v_{2}+v_{3}+v_{4}-v_{5}-v_{6}-v_{7}-v_{8}+u_{1}+u_{2}+u_{3}+u_{4}-u_{5}-u_{6}$, now we calculate the upper bound of $I$.
 Firstly, regarding $x\leqslant 2U^{4}$ or $x\geqslant 32U^{4}$, in both cases, the integral region does not exist, so we have $2U^{4}<x<32U^{4}$.
-In this case, we record as $u_{3}=xu$, then there is
+In this case, we record as $u_{7}=xu$, then there is
 $\begin{split}I&\leqslant(1+\varepsilon)x^{-1/2}L^{-1}\int_{\max(U^{4}/x,1-16U^{4}/x)}^{\min(16U^{4}/x,1-U^{4}/x)}u^{-3/4}(1-u)^{-3/4}\,du\\ &\leqslant\frac{1+\varepsilon}{\sqrt{2}}U^{-2}L^{-1}\int_{1/17}^{16/17}u^{-3/4}(1-u)^{-3/4}\,du\\ &\leqslant\frac{1+\varepsilon}{\sqrt{2}}U^{-2}L^{-1}I^{*},\end{split}$
 where we record as $I^{*}=\int_{1/17}^{16/17}u^{-3/4}(1-u)^{-3/4}\,du$, then substituting it into (3.3) yields the following
-$J(n)$ $\leq\frac{1+\varepsilon}{4\sqrt{2}}I^{*}U^{-2}L^{-1}\Bigg(\>\int_{U^{4}}^{16U^{4}}\>\frac{du}{u^{3/4}\log u}\Bigg)^{2}\Bigg(\>\int_{V^{4}}^{16V^{4}}\>\frac{dv}{v^{3/4}\log v}\Bigg)^{4}+O(V^{4}L^{-3B})$
-$<\left(\frac{24}{5}\right)^{4}\frac{4}{\sqrt{2}}(1+\varepsilon)I^{*}V^{4}L^{-7}+O(V^{4}L^{-3B})$
-$<11606V^{4}L^{-7},$
-where we have used the estimate $I^{*}<7.73$. This proves (ii). For the proofs of (i) and (iii), we can refer to the methods of Brüdern [1] and Ren[2]. This proves Lemma 1.
+$J(n)$ $\leq\frac{1+\varepsilon}{4\sqrt{2}}I^{*}U^{-2}L^{-1}\Bigg(\>\int_{U^{4}}^{16U^{4}}\>\frac{du}{u^{3/4}\log u}\Bigg)^{6}\Bigg(\>\int_{V^{4}}^{16V^{4}}\>\frac{dv}{v^{3/4}\log v}\Bigg)^{8}+O(U^{4}V^{8}L^{-5B})$
+$<\left(\frac{32}{7}\right)^{8}\frac{4^{5}}{\sqrt{2}}(1+\varepsilon)I^{*}U^{4}V^{8}L^{-15}+O(U^{4}V^{8}L^{-5B})$
+$<4888799.222U^{4}V^{8}L^{-15},$
+where we have used the estimate $I^{*}<7.73$. This proves (ii). Now we prove (iii), let
+$v_{d}(n)=\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0({\mathrm{mod}}~d)\end{subarray}}R(m);f_{d}(\alpha)=\sum_{U<x\leq 2U\atop x\equiv 0\mod d}e(\alpha x^{4});$
+$g(\alpha)=\sum_{U<p\leq 2U}e(\alpha p^{4});h(\alpha)=\sum_{V<p\leq 2V}e(\alpha p^{4})$
+and then define
+$F(\alpha)=\sum_{d<D}\eta_{d}f_{d}(\alpha),$
+$v_{d}(n,\mathfrak{B})=\int_{\mathfrak{B}}f_{d}(\alpha)\overline{g(\alpha)}|g(\alpha)|^{4}|h(\alpha)|^{8}e(-\alpha n)d\alpha.$
+Naturally, $v_{d}(n,[0,1])=v_{d}(n)$, therefore
+(3.4) $\begin{split}|\sum_{d\leqslant D}\eta_{d}E_{d}(n)|&\leq\sum_{d\leqslant D}|\eta_{d}||v_{d}(n)-\frac{\mathfrak{S}_{d}(n)}{d}J(n)|\\ &\leq\sum_{d\leqslant D}|\eta_{d}||v_{d}(n,\mathfrak{M})-\frac{\mathfrak{S}_{d}(n)}{d}J(n)|+\sum_{d\leqslant D}|\eta_{d}||v_{d}(n,\mathfrak{m})|.\end{split}$
+Next, we calculate the second part on the right side of the inequality
+(3.5) $\begin{split}\sum_{d\leqslant D}|\eta_{d}||v_{d}(n,\mathfrak{m})|&=\int_{\mathfrak{m}}|F(\alpha)||g(\alpha)|^{7}|h(\alpha)|^{8}e(-\alpha n)d\alpha\\ &\leq\bigg(\int_{\mathfrak{m}}|F(\alpha)|^{2}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha\bigg)^{1/2}\bigg(\int_{\mathfrak{m}}|g(\alpha)|^{12}|h(\alpha)|^{12}d\alpha\bigg)^{1/2}.\end{split}$
+Now we calculate the upper bound of $F(\alpha)$. Regarding the above $\mathfrak{M}^{\prime}(q,a)$, $1\leq a\leq q\leq U^{2}$, we write $F(\alpha)$ as
+$F(\alpha)=\sum_{d\leq D}\eta_{d}\sum_{U/d<y\leq 2U/d}e(\alpha d^{4}y^{4}).$
+By Dirichlet’ s theorem on diophantine approximation, there are coprime integers $b$, $r$ with $r\leq 16P^{3}d^{-3}$, $\left|d^{4}\alpha-\frac{b}{r}\right|\leq\frac{1}{16}r^{-1}d^{3}P^{-3}$.
+By Weyl’ s inequality, when $r>U/d$
+$\sum_{U/d<y\leq 2U/d}e(\alpha d^{4}y^{4})\ll\bigg(\frac{U}{d}\bigg)^{7/8},$
+and when $r\leq U/d$
+(3.6) $\sum_{U/d<y\leq 2U/d}e(\alpha d^{4}y^{4})\ll r^{-\frac{1}{4}}\frac{U}{d}\bigg(1+\bigg(\frac{U}{d}\bigg)^{4}\bigg|\alpha d^{4}-\frac{b}{r}\bigg|\bigg)^{-\frac{1}{4}}+\bigg(\frac{P}{d}\bigg)^{\frac{1}{2}+\epsilon}.$
+Furthermore, when
+$r\leq(U/d)^{7/8};\quad\left|\alpha d^{4}-\frac{b}{r}\right|\leq\frac{1}{r}\left(\frac{d}{U}\right)^{\frac{7}{2}},$
+it can also achieve a result of (3.6). Hence
+$F(\alpha)\ll U^{7/8+\epsilon}D^{1/8}+U\sum_{d\in\mathcal{D}}d^{-1}r^{-\frac{1}{4}}\biggl(1+\biggl(\frac{U}{d}\biggr)^{4}\biggl|\alpha d^{4}-\frac{b}{r}\biggr|\biggr)^{-\frac{1}{4}},$
+where $\mathcal{D}$ represents the set of d that satisfies the condition. Compare the conditions of $q,a$ and $b,r$, we have
+$\bigg|\frac{b}{r}-\frac{ad^{4}}{q}\bigg|\leq\frac{1}{r}\bigg(\frac{d}{U}\bigg)^{\frac{7}{2}}+\frac{d^{4}}{qU^{2}}$
+i.e.
+$\bigg|bq-ad^{4}r\bigg|\leq qd^{\frac{7}{2}}U^{-\frac{7}{2}}+rU^{-2}D^{4}\ll 1$
+where $D=N^{\xi}$, $0<\xi<9/25$, $U$ is large enough. Therefore $bq=ad^{4}r$, then $r=q/(q,d^{4})$, by the trivial bound $(q,d^{4})\leq(q,d)^{4}$
+$\sum_{d\in\mathcal{D}}d^{-1}r^{-\frac{1}{4}}\left(1+\left(\frac{U}{a}\right)^{4}\left|\alpha d^{4}-\frac{b}{r}\right|\right)^{-\frac{1}{4}}\leq q^{-\frac{1}{4}}\left(1+P^{4}\left|\alpha-\frac{a}{q}\right|\right)^{-\frac{1}{4}}\sum_{d\leq D}\frac{(q,d)}{d}.$
+Thus
+$F(\alpha)\ll U^{\frac{7}{8}+\epsilon}D^{\frac{1}{8}}+q^{\epsilon-\frac{1}{4}}P(\log P)\left(1+P^{4}\bigg|\alpha-\frac{a}{q}\bigg|\right)^{-\frac{1}{4}}.$
+Let $\delta>0$ be so small that $D^{\frac{1}{8}}U^{\frac{7}{8}}\leq U^{\frac{23}{25}-2\delta}$, this is always possible because $D<U^{\frac{9}{25}}$.
+Let $\mathfrak{N}(q,a)$ denote the interval $|q\alpha-a|\leqslant U^{5\delta-\frac{92}{25}}$, and $\mathfrak{N}$ be the union of all $\mathfrak{N}(q,a)$ with $1\leqslant a\leqslant q\leqslant P^{\frac{8}{25}+5\delta}$, $(a,q)=1$.
+By the upper bound of $F(\alpha)$ obtained from the above, for $|F(\alpha)>U^{\frac{23}{25}}|$, then $\alpha\in\mathfrak{N}$.
+Defining $\Phi$ on $\mathfrak{N}$ as
+$\Phi(\alpha)=q^{\epsilon-\frac{1}{4}}\Bigg(1+U^{4}\Bigg|\alpha-\frac{a}{q}\Bigg|\Bigg)^{-1/4}$
+Then we have
+(3.7) $\int_{\mathfrak{m}}|Fgh^{2}|^{2}d\alpha\ll U^{\frac{46}{25}-2\delta}\int_{0}^{1}|gh^{2}|^{2}d\alpha+U^{2}(\log U)^{2}\int_{\mathfrak{N}\cap\mathfrak{m}}|\Phi gh^{2}|^{2}d\alpha.$
+By Hölder’ s inequality
+(3.8) $\int_{\mathfrak{N}\cap\mathfrak{m}}|\Phi gh^{2}|^{2}d\alpha\leq\left(\int_{\mathfrak{N}\cap\mathfrak{m}}|\Phi|^{12}d\alpha\right)^{\frac{1}{6}}U_{1}^{\frac{1}{2}}U_{2}^{\frac{1}{3}}$
+where
+$U_{1}=\int_{0}^{1}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha;\quad U_{2}=\int_{0}^{1}|g(\alpha)|^{3}|h(\alpha)|^{6}d\alpha.$
+By simple calculations we obtain that
+$\int_{\mathfrak{N}\cap\mathfrak{m}}|\Phi(\alpha)|^{12}d\alpha\ll U^{-4}L^{-B}.$
+Then let $\mathfrak{m}_{1}(q,a)$ denote the interval $|q\alpha-a|\leqslant U^{\frac{12}{5}}$, $\mathfrak{m}_{1}$ be the union of all $\mathfrak{m}_{1}(q,a)$ with $1\leqslant a\leqslant q\leqslant U^{\frac{4}{5}}$, $(a,q)=1$, and $\mathfrak{m}_{2}(q,a)$ denote the interval $|q\alpha-a|\leqslant V^{\frac{12}{5}}$, $\mathfrak{m}_{2}$ be the union of all $\mathfrak{m}_{2}(q,a)$ with $1\leqslant a\leqslant q\leqslant V^{\frac{4}{5}}$, $(a,q)=1$.
+By Theorem 4.1, Lemma 6.3 of Vaughan[7], and Theorem 2 of Vaughan[11], we have
+$g(\alpha)\ll q^{-\frac{1}{4}}U\bigg(1+U^{4}\bigg|\alpha-\frac{a}{q}\bigg|\bigg)^{-1}+U^{4/5+\epsilon}$
+where $\alpha\in\mathfrak{m}_{1}(q,a)$;
+$h(\alpha)\ll q^{-\frac{1}{4}}V\bigg(1+V^{4}\bigg|\alpha-\frac{a}{q}\bigg|\bigg)^{-1}+V^{4/5+\epsilon}$
+where $\alpha\in\mathfrak{m}_{2}(q,a).$ Moreover, $|g(\alpha)|>U^{4/5+2\epsilon}$ implies $\alpha\in{\mathfrak{m}_{1}}({\mathrm{mod}}~1)$, $|h(\alpha)|>V^{4/5+2\epsilon}$ implies $\alpha\in{\mathfrak{m}_{2}}({\mathrm{mod}}~1)$.
+(3.9) $g(\alpha)\ll q^{-1/4}U+U^{1/2}\ll q^{-1/4}U$
+where $\alpha\in{\mathfrak{m}_{1}}({\mathrm{mod}}~1)$;
+(3.10) $h(\alpha)\ll q^{-1/4}V+V^{1/2}\ll q^{-1/4}V$
+where $\alpha\in{\mathfrak{m}_{2}}({\mathrm{mod}}~1).$ Then define $\Psi(\alpha)$ on $\mathfrak{m}_{1}$, $\Psi^{*}(\alpha)$ on $\mathfrak{m}_{2}$ as
+(3.11) $\Psi(\alpha)=q^{-\frac{1}{4}}\left(1+U^{4}\bigg|\alpha-\frac{a}{q}\bigg|\right)^{-1};\Psi^{*}(\alpha)=q^{-\frac{1}{4}}\left(1+V^{4}\bigg|\alpha-\frac{a}{q}\bigg|\right)^{-1},$
+Subsequently
+$\int_{0}^{1}|g(\alpha)|^{3}|h(\alpha)|^{6}d\alpha\ll$ $\>U^{\frac{4}{5}}\int_{0}^{1}|g(\alpha)|^{2}|h(\alpha)|^{6}d\alpha+U\int_{\mathfrak{m}_{1}}|\Psi(\alpha)|^{2}|g(\alpha)|^{2}|h(\alpha)|^{6}d\alpha$
+$\ll$ $\>U^{\frac{4}{5}}V^{\frac{8}{5}}\int_{0}^{1}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha+U^{\frac{4}{5}}V^{2}\int_{\mathfrak{m}_{2}}|\Psi^{*}(\alpha)|^{2}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha$
+$+UV^{\frac{8}{5}}\int_{\mathfrak{m}_{1}}|\Psi(\alpha)|^{2}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha$
+$+UV^{2}\int_{\mathfrak{m}_{1}\cap\mathfrak{m}_{2}}|\Psi(\alpha)|^{2}|\Psi^{*}(\alpha)|^{2}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha.$
+By (3.9), (3.10), (3.11) we can obtain that
+(3.12) $U_{2}=\int_{0}^{1}|g(\alpha)|^{3}|h(\alpha)|^{6}d\alpha\ll U^{5+13/40}.$
+Since the upper bound of $U_{1}$ is known, we can combine(3.7), (3.8), (3.12) to obtain that
+$\int_{\mathfrak{m}}|Fgh^{2}|^{2}d\alpha\ll U^{4+193/200}L^{-B/6}.$
+Similar to the method mentioned above, we can obtain
+$\int_{\mathfrak{m}}|g(\alpha)|^{12}|h(\alpha)|^{12}d\alpha\ll U^{16+29/40}$
+Naturally, by (3.5)
+$\sum_{d\leqslant D}|\eta_{d}||v_{d}(n,\mathfrak{m})|\ll U^{11}L^{-B/12}$
+Now we calculate the first term on the right side of the inequality (3.4). For $\alpha=a/q+\beta\in\mathfrak{M}$, we define
+$f_{d}^{*}(\alpha)=\frac{S(q,ad^{4})}{qd}\Phi(\lambda,U),\quad g^{*}(\alpha)=\frac{C(q,a)}{\varphi(q)}\Psi(\lambda,U),\quad h^{*}(\alpha)=\frac{C(q,a)}{\varphi(q)}\Psi(\lambda,V).$
+By Lemma 7.15 of Hua[12]
+$g(\alpha)-g^{*}(\alpha)\ll U\exp(-c_{1}\sqrt{L}),\quad h(\alpha)-h^{*}(\alpha)\ll V\exp(-c_{1}\sqrt{L})$
+where $c_{1}>0$ is a absolute constant.By Theorem 4.1 of Vaughan[11]
+$f_{d}(\alpha)-f_{d}^{*}(\alpha)\ll q^{1/2+\epsilon}(1+U^{4}|\lambda|)\ll L^{2B}.$
+Then for $\alpha\in\mathfrak{M}$, by trivial estimate $\left|S(q,ad^{3})\right|\leq q$, we have
+$f_{d}\overline{g}|g|^{6}|h|^{8}-f_{d}^{*}\overline{g^{*}}|g^{*}|^{6}|h^{*}|^{8}\ll d^{-1}U^{8}V^{8}L^{-13B},$
+and
+(3.13) $\sum_{d\leq D}|\eta_{d}||\nu_{d}(n,\mathfrak{M})-\nu_{d}^{*}(n)|\ll U^{4}V^{8}L^{-A},$
+where
+$\nu_{d}^{*}(n)=\int_{\mathfrak{M}}f_{d}^{*}\overline{{g^{*}}}|g^{*}|^{6}|h^{*}|^{8}e(-n\alpha)d\alpha,$
+and we can find that
+(3.14) $\nu_{d}^{*}(n)=\frac{1}{d}J(n)\sum_{q\leq L^{B}}T_{d}(n,q).$
+Then by $|C(q,a)|\ll q^{3/4+\epsilon}$
+$|T_{d}(n,q)|\leq\sum_{a=1}^{q}\frac{|S(q,ad^{4})||C(q,a)|^{15}}{q\varphi^{15}(q)}\ll\sum_{a=1}^{q}\frac{q^{49/4+\epsilon}}{q\varphi^{15}(q)}\ll q^{-2}.$
+Hence the series
+$\mathfrak{S}_{d}(n)=\sum_{q=1}^{\infty}T_{d}(n,q)$
+converges absolutely, this proves (i), and
+$\sum_{q\leq L^{B}}T_{d}(n,q)=\mathfrak{S}_{d}(n)+O(L^{-B}).$
+By (3.14)
+$\nu_{d}^{*}(n)=\frac{\mathfrak{S}_{d}(n)}{d}J(n)+O\left(\frac{U^{4}V^{8}}{dL^{B}}\right),$
+combining with (3.13), we can obtain
+$\sum_{d\leqslant D}|\eta_{d}||v_{d}(n,\mathfrak{M})-\frac{\mathfrak{S}_{d}(n)}{d}J(n)|\ll U^{4}V^{8}L^{-A}.$
+This proves (iii), which also proves Lemma 2.
 Lemma 3. For $(d,6)=1$, we get
@@
 Lemma 4. Define $K(n,p)$ as the number of solutions to the following equation
-$y_{2}^{4}+\cdots+y_{4}^{4}-y_{5}^{4}-\cdots-y_{8}^{4}\equiv n(\text{mod}~p)$
-where $1\leq y_{i}\leq p-1$ $(1\leq i\leq 8)$, then we get
-$pK(n,p)=(p-1)^{7}+E,$
+$y_{2}^{4}+\cdots+y_{8}^{4}-y_{9}^{4}-\cdots-y_{16}^{4}\equiv n(\text{mod}~p)$
+where $1\leq y_{i}\leq p-1$ $(2\leq i\leq 16)$, then we get
+$pK(n,p)=(p-1)^{15}+E,$
 where
@@
 and
-$|E|\leqslant(3\sqrt{p}+1)^{5}(p-1)(3p+1),~~~\textit{if}~p\equiv 1({\mathrm{mod}}~4).$
+$|E|\leqslant(3\sqrt{p}+1)^{{13}}(p-1)(3p+1),~~~\textit{if}~p\equiv 1({\mathrm{mod}}~4).$
 It follows that $K(n,p)>0$ for $p\geq 17$.
 Proof. From the definition of $K(n,p)$, we have
-$pK(n,p)=\sum_{a=1}^{p}C^{3}(p,a)\overline{C^{4}(p,a)}e\left(-\frac{an}{p}\right)=(p-1)^{7}+E,$
+$pK(n,p)=\sum_{a=1}^{p}C^{7}(p,a)\overline{C^{8}(p,a)}e\left(-\frac{an}{p}\right)=(p-1)^{15}+E,$
 where
-$E=\sum_{a=1}^{p-1}C^{3}(p,a)\overline{C^{4}(p,a)}e\biggl(-\frac{an}{p}\biggr).$
+$E=\sum_{a=1}^{p-1}C^{7}(p,a)\overline{C^{8}(p,a)}e\biggl(-\frac{an}{p}\biggr).$
 When $p\equiv 3(\text{mod }4)$ and $(p,a)=1$, by Lemma 4.3 in Vaughan[7], we get $S(p,a)=0$ thus C(p,a)=1C(p,a)=-1, therefore
@@
 $\sum_{a=1}^{p-1}|C(p,a)|^{2}=\sum_{a=1}^{p}|C(p,a)|^{2}-(p-1)^{2}.$
-Obiviously, $\sum_{a=1}^{p}|C(p,a)|^{2}$ can be expressed as $p$ times the number of solutions to equation $x^{4}\equiv y^{4}({\mathrm{mod~}}p)$, $1\leqslant x$, $y\leqslant p-1$.
+Obviously, $\sum_{a=1}^{p}|C(p,a)|^{2}$ can be expressed as $p$ times the number of solutions to equation $x^{4}\equiv y^{4}({\mathrm{mod~}}p)$, $1\leqslant x$, $y\leqslant p-1$.
 For $p\equiv 1({\mathrm{mod~}}4)$,
@@
 thus
-$\sum_{a=1}^{p-1}|C^{3}(p,a)\overline{C^{4}(p,a)}|\leqslant(3\sqrt{p}+1)^{5}(p-1)(3p+1).$
+$\sum_{a=1}^{p-1}|C^{7}(p,a)\overline{C^{8}(p,a)}|\leqslant(3\sqrt{p}+1)^{13}(p-1)(3p+1).$
 A simple calculation shows that $K(n,p)>0$ for $p\geq 15$, $p\equiv 1(\text{mod }4)$and for all $p\equiv 3(\text{mod }4)$, therefore we have chosen $p\geq 17$. This proves Lemma 4.
@@
 Lemma 5. For $i=1,2$, let $H(n,p^{i})$ denote the number of solution of
-$x^{3}+y_{2}^{3}+\cdots+y_{4}^{3}-y_{5}^{3}-\cdots-y_{8}^{3}\equiv n\pmod{p^{i}}$
+$x^{4}+y_{2}^{4}+\cdots+y_{8}^{4}-y_{9}^{4}-\cdots-y_{16}^{4}\equiv n\pmod{p^{i}}$
 where $1\leq x\leq p^{i},1\leq y_{j}<p^{i}$ and $(y_{j},p)=1$. Thus
-$pH(n,p)=p(p-1)^{7}+E^{*},$
+$pH(n,p)=p(p-1)^{15}+E^{*},$
 where
@@
 and
-$|E^{*}|\leq 3\sqrt{p}(3\sqrt{p}+1)^{5}(p-1)(3p+1),~~\textit{if}~p\equiv 1(\text{mod }4).$
-It follows that $H(n,p)>0$ for all pp.
+$|E^{*}|\leq 3\sqrt{p}(3\sqrt{p}+1)^{13}(p-1)(3p+1),~~\textit{if}~p\equiv 1(\text{mod }4).$
+It also follows that $H(n,p)>0$ for $p>17$.
 Lemma 6. Definne the functions
@@
 Lemma 7. (Mertens’ theorem)For prime number $p$, $x\geq e$, positive integers $k$, $l$ with $(k,l)=1$ and $k\leq\ln^{A}{x}$ for any positive integer A, then
-$\prod_{p\leq x,p\equiv l({\mathrm{mod~}k})}\left(1-\frac{1}{p}\right)=\frac{e^{-\frac{1}{\varphi(k)}\left[\gamma+\ln\frac{\varphi(k)}{k}\right]}}{\ln^{\frac{1}{\varphi(k)}}x}\left\{1+O\left(e^{-c\ln^{\frac{3}{5}}x}\right)\right\}$
-where $c$ is a positive absolute constant, $\gamma$ is the Euler’s constant, and the constant in $O$ is independent of $x$.
+$\prod_{\begin{subarray}{c}p\leq x\\ p\equiv l({\mathrm{mod~}k})\end{subarray}}\left(1-\frac{1}{p}\right)=\frac{e^{-\frac{1}{\varphi(k)}\left[\gamma+\ln\frac{\varphi(k)}{k}\right]}}{\ln^{\frac{1}{\varphi(k)}}x}\left\{1+O\left(e^{-c\ln^{\frac{3}{5}}x}\right)\right\}$
+where $c$ is a positive absolute constant, $\gamma$ is the Euler’s constant, and the constant in $O$ is independent of $x$. Proof. See [15], corollary of Theorem 429.
 Lemma 8. For $N/9<\ell\leqslant N$, define $r(\ell)$ as the number of $\ell$ can be expressed in the form of (1.2) with
-$p_{1},p_{2}\sim U,\qquad p_{3},p_{4}\sim V\,.$
+$p_{1},p_{2},p_{3},p_{4}\sim U,\qquad p_{5},p_{6},p_{7},p_{8}\sim V\,.$
 Then we have
-$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)<bV^{4}L^{-8},$
-where $b=3702014554$.
-Proof. Denote $\mathcal{B}$ as the set of all prime numbers greater than 13, and define
+$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)<bU^{2}V^{8}L^{-14},$
+where $b=80947432211.141$.
+Proof. Denote $\mathcal{B}$ as the set of all prime numbers greater than 17, and define
 $P(z)=\prod_{\begin{subarray}{c}p<z\\ p\in\mathcal{B}\end{subarray}}p.$
 To prove Lemma 8, we hope to obtain an appropriate upper bound for the following equation
-$\sum_{m\sim U\atop(m,P(z))=1}R(m).$
+$\sum_{\begin{subarray}{c}m\sim U\\ \atop(m,P(z))=1\end{subarray}}R(m).$
 For this, we apply Lemma 6, there is no prime divisor beyond $\mathcal{B}$ for $d$, then we define
@@
 By (2.3), we get
-$1+T_{p}(n,p)=\sum_{a=1}^{p}\frac{\overline{C(p,a)}|C(p,a)|^{6}}{(p-1)^{7}}e\Bigg(-\frac{an}{p}\Bigg)=p\frac{K(n,p)}{(p-1)^{7}}.$
+$1+T_{p}(n,p)=\sum_{a=1}^{p}\frac{\overline{C(p,a)}|C(p,a)|^{14}}{(p-1)^{15}}e\Bigg(-\frac{an}{p}\Bigg)=p\frac{K(n,p)}{(p-1)^{15}}.$
 hence
@@
 Now let $X=\mathfrak{S}_{1}(n)J(n)$, then by Lemma 2, we have
-$\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0\mod d\end{subarray}}R(m)=\frac{\omega(d)}{d}X+E_{d}(n).$
-Suppose $U_{0}\geqslant 1,\ V_{0}\geqslant 1,\ U_{0}V_{0}=D=N^{(1-\epsilon)/16}$ and $z=D^{1/2}$, then
+$\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0(\mod d)\end{subarray}}R(m)=\frac{\omega(d)}{d}X+E_{d}(n).$
+Suppose $U_{0}\geqslant 1,\ V_{0}\geqslant 1,\ U_{0}V_{0}=D=N^{9(1-\epsilon)/100}$ and $z=D^{1/2}$, then
 $\frac{\log(U_{0}V_{0})}{\log z}=2.$
@@
 by Lemma 2, we have
-$\sum_{1\leqslant m\leqslant U}a_{m}\sum_{1\leqslant k\leqslant V}b_{k}E_{mk}\ll\sum_{1\leqslant d\leqslant D}\tau(d)E_{d}\ll V^{4}L^{-A}.$
-By Lemma 5, we get
+$\sum_{1\leqslant m\leqslant U}a_{m}\sum_{1\leqslant k\leqslant V}b_{k}E_{mk}\ll\sum_{1\leqslant d\leqslant D}\tau(d)E_{d}\ll U^{4}V^{8}L^{-A}.$
+By the upper bound in Lemma 6, we get
 $\sum\limits_{(m,P(z))=1}R(m)<e^{\gamma}(1+\varepsilon)J(n)\mathfrak{S}_{1}(n)W(z),$
@@
 Actually, we have estimated the upper bound of $J(n)$ in Lemma 2. Next, we estimate the remaining part,
-(3.2) $\begin{split}&\mathfrak{S}_{1}(n)W(z)\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))\\ &\times\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}(1+T_{1}(n,p))\biggl(1-\frac{K(n,p)}{H(n,p)}\biggr)\\ &\times\prod_{p\geqslant N^{(1-\varepsilon)/32}}(1+T_{1}(n,p))\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))\\ &\times\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1-\frac{1}{p}\right)\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\\ &\times\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}}{p(p-1)^{7}}\right).\end{split}$
+(3.15) $\begin{split}&\mathfrak{S}_{1}(n)W(z)\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))(1+T_{1}(n,13))\\ &\times\prod_{17\leqslant p<N^{9(1-\varepsilon)/200}}(1+T_{1}(n,p))\biggl(1-\frac{K(n,p)}{H(n,p)}\biggr)\\ &\times\prod_{p\geqslant N^{9(1-\varepsilon)/200}}(1+T_{1}(n,p))\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))(1+T_{1}(n,13))\\ &\times\prod_{17\leqslant p<N^{(1-\varepsilon)/200}}\left(1-\frac{1}{p}\right)\left(1+\frac{E^{*}-E}{\left(p-1\right)^{16}}\right)\\ &\times\prod_{p\geqslant N^{(1-\varepsilon)/200}}\left(1+\frac{E^{*}}{p(p-1)^{15}}\right).\end{split}$
 For $n$ in the definition of $R(m)$, define $\rho(n)$ as the number of solutions to the following equation
-$n=p_{1}^{4}+\cdots+p_{4}^{4}-p_{5}^{4}-\cdots-p_{8}^{4},$
+$n=p_{1}^{4}+\cdots+p_{8}^{4}-p_{9}^{4}-\cdots-p_{16}^{4},$
 where
-$p_{1},p_{2},p_{5},p_{6}\sim U,\quad p_{3},p_{4},p_{7},p_{8}\sim V.$
+$p_{1},p_{2},p_{3},p_{4},p_{9},p_{10},p_{11},p_{12}\sim U,\quad p_{5},p_{6},p_{7},p_{8},p_{13},p_{14},p_{15},p_{16}\sim V.$
 We can naturally obtain
@@
 According to the definition (2.3), (2.4), by simple calculation we have $S(2,1)=0$, then
-$T_{1}(n,2)=\sum_{\overset{a=1}{(a,2)=1}}^{2}\frac{S(2,a)C^{3}(2,a)\overline{C^{4}(2,a)}}{2\varphi^{7}(2)}e\biggl(-\frac{an}{2}\biggr)=0.$
+$T_{1}(n,2)=\sum_{\overset{a=1}{(a,2)=1}}^{2}\frac{S(2,a)C^{7}(2,a)\overline{C^{8}(2,a)}}{2\varphi^{16}(2)}e\biggl(-\frac{an}{2}\biggr)=0.$
 For $q=4$, we have
-$T_{1}(n,4)=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{S(4,a)C^{3}(4,a)\overline{C^{4}(4,a)}}{4\varphi^{7}(4)}e\biggl(-\frac{an}{4}\biggr),$
+$T_{1}(n,4)=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{S(4,a)C^{7}(4,a)\overline{C^{8}(4,a)}}{4\varphi^{16}(4)}e\biggl(-\frac{an}{4}\biggr),$
 where
@@
 Thus
-$\begin{split}T_{1}(n,4)&=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{(2+2e(\frac{a}{4}))(2e(\frac{a}{4}))^{3}\overline{2e(\frac{a}{4})}^{4}}{4\varphi^{7}(4)}e\biggl(-\frac{a\times 0}{4}\biggr)\\ &=\frac{1}{2}\sum_{\overset{a=1}{(a,4)=1}}^{4}(1+e(-\frac{a}{4}))=\frac{1}{2}(2+2\cos{\frac{\pi}{2}})=1.\end{split}$
+$\begin{split}T_{1}(n,4)&=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{(2+2e(\frac{a}{4}))(2e(\frac{a}{4}))^{7}\overline{2e(\frac{a}{4})}^{8}}{4\varphi^{13}(4)}e\biggl(-\frac{a\times 0}{4}\biggr)\\ &=\frac{S(4,1)C^{7}(4,1)+S(4,3)C^{7}(4,3)}{2^{15}}=1.\end{split}$
 Similarly, we have
-$T_{1}(n,8)=\frac{1}{2}(4+2\cos{\frac{\pi}{4}}+2\cos{\frac{3\pi}{4}})=2$
+$T_{1}(n,8)=\sum_{\overset{a=1}{(a,8)=1}}^{8}\frac{1}{2}\bigg(1+e\bigg(\frac{7a}{8}\bigg)\bigg)=2$
 and
-$T_{1}(n,16)=\frac{1}{2}(8+2\cos{\frac{\pi}{8}}+2\cos{\frac{3\pi}{8}}++2\cos{\frac{5\pi}{8}}+2\cos{\frac{7\pi}{8}})=4.$
+$T_{1}(n,16)=\sum_{\overset{a=1}{(a,16)=1}}^{16}\frac{1}{2}\bigg(1+e\bigg(\frac{15a}{16}\bigg)\bigg)=4.$
 Thus
@@
 Also, for $q=3$, we have
-$T_{1}(n,3)=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{S(3,a)C^{3}(3,a)\overline{C^{4}(3,a)}}{3\varphi^{7}(3)}e\biggl(-\frac{an}{3}\biggr),$
+$T_{1}(n,3)=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{S(3,a)C^{7}(3,a)\overline{C^{8}(3,a)}}{3\varphi^{15}(3)}e\biggl(-\frac{an}{3}\biggr),$
 where
-$S(3,a)=\sum_{m=1}^{3}e\biggl(\frac{am^{4}}{3}\biggr)=1+2e(\frac{a}{3}),C(4,a)=\sum_{(m,4)=1}^{4}e\biggl(\frac{am^{4}}{4}\biggr)=2e(\frac{a}{3}).$
+$S(3,a)=\sum_{m=1}^{3}e\biggl(\frac{am^{4}}{3}\biggr)=1+2e\bigg(\frac{a}{3}\bigg),C(4,a)=\sum_{(m,4)=1}^{4}e\biggl(\frac{am^{4}}{4}\biggr)=2e\bigg(\frac{a}{3}\bigg).$
 then
-$\begin{split}T_{1}(n,3)&=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{(1+2e(\frac{a}{3}))(2e(\frac{a}{3}))^{3}\overline{(2e(\frac{a}{3}))^{4}}}{3\varphi^{7}(3)}e\biggl(-\frac{a\times 0}{3}\biggr)\\ &=\frac{1}{3}(4+e(-\frac{1}{3})+e(-\frac{2}{3}))=\frac{1}{3}(4+2\cos{\frac{2\pi}{3}})=1.\end{split}$
-Likewise, by a series of calculations we have $T_{1}(n,5)=\frac{1}{5}(16+2\cos{\frac{2\pi}{5}}+2\cos{\frac{4\pi}{5}})=\frac{1}{5}(16+\frac{\sqrt{5}-1}{2}-\frac{\sqrt{5}+1}{2})=3$, $T_{1}(n,7)=\frac{8}{729}$ and $T_{1}(n,11)=\frac{324}{171875}.$
+$\begin{split}T_{1}(n,3)&=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{(1+2e(\frac{a}{3}))(2e(\frac{a}{3}))^{7}\overline{(2e(\frac{a}{3}))^{8}}}{3\varphi^{15}(3)}e\biggl(-\frac{a\times 0}{3}\biggr)=1.\end{split}$
+Likewise, by a series of calculations we have $T_{1}(n,5)=3$, $T_{1}(n,7)=\frac{128}{3^{14}}$, $T_{1}(n,11)=\frac{2187}{6103515625}$ and $T_{1}(n,13)\leq 0.015$ .
 Furthermore, for
-$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right),$
+$\prod_{17\leqslant p<N^{9(1-\varepsilon)/200}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{16}}\right),$
 by Lemma 4 and Lemma 5, for $p\equiv 3(\text{mod }4)$ we have
@@
 then
-$\begin{split}\prod_{\overset{13\leqslant p<N^{(1-\varepsilon)/32}}{p\equiv 3(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)&\leq\prod_{\overset{p\geqslant 13}{p\equiv 3(\mathrm{mod~4})}}\left(1+\frac{1}{\left(p-1\right)^{7}}\right)\\ &<1\>+\>\sum_{n\geqslant 12}\>\frac{1}{n^{7}}<1\>+\>12^{-6},\end{split}$
+$\begin{split}\prod_{\overset{17\leqslant p<N^{9(1-\varepsilon)/200}}{p\equiv 3(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{16}}\right)&\leq\prod_{\overset{p\geqslant 17}{p\equiv 3(\mathrm{mod~4})}}\left(1+\frac{1}{\left(p-1\right)^{15}}\right)\\ &<1\>+\>\sum_{n\geqslant 19}\>\frac{1}{n^{15}}<1\>+\>19^{-14},\end{split}$
 and for $p\equiv 1(\text{mod }4)$, we have
-$|E-E^{*}|\leq(3\sqrt{p}+1)^{6}(p-1)(3p+1).$
-then for $p\geq 101$, by $1+x\leq e^{x}$, we have
-$\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}e^{\frac{E^{*}-E}{\left(p-1\right)^{8}}}=e^{\sum_{p\geq 101}{\frac{E^{*}-E}{\left(p-1\right)^{8}}}},$
-where
-$\sum_{p\geq 101}{\frac{|E^{*}-E|}{\left(p-1\right)^{8}}}\leq 0.305,$
-thus
-$\prod_{\overset{p\geq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq 1.357.$
-In addition,
-$\prod_{\overset{13\leq p\leq 101}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\leq 11.773.$
+$|E-E^{*}|\leq(3\sqrt{p}+1)^{14}(p-1)(3p+1).$
+Now define $\delta_{p}$ as
+$\left|\frac{E^{*}-E}{(p-1)^{16}}\right|\leq\frac{(3\sqrt{p}+1)^{14}(3p+1)}{(p-1)^{15}}=\delta_{p}.$
+By numerical calculation
+$\sum_{p\geq 17}\delta_{p}<0.03,$
+then by $1+x\leq e^{x}$
+$\prod_{\overset{17\leq p\leq N^{{9(1-\epsilon)}/{200}}}{p\equiv 1(\text{mod}4)}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{16}}\right)\leq\prod_{\overset{p\geq 17}{p\equiv 1(\text{mod}4)}}e^{\frac{E^{*}-E}{\left(p-1\right)^{16}}}\leq e^{\sum_{p\geq 17}{\delta_{p}}}\leq 1.1.$
 Therefore we have
-$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)<15.976$
+$\prod_{13\leqslant p<N^{(9(1-\varepsilon)/200}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{16}}\right)<1.1$
 Furthermore, by Lemma 7, we have the estimate
-$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1-\frac{1}{p}\right)<308e^{-\gamma}(1+\varepsilon)L^{-1}.$
-It is not difficult to find that
-$\frac{|E^{*}|}{p(p-1)^{7}}<\frac{1}{p^{2}}$
-for $p$ large enough. Hence
-$\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}}{p(p-1)^{7}}\right)\leqslant\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{1}{p^{2}}\right)<1+\varepsilon.$
+$\prod_{17\leqslant p<N^{(1-\varepsilon)/32}}\left(1-\frac{1}{p}\right)<231.713e^{-\gamma}(1+\varepsilon)L^{-1}.$
+It is not difficult to find that for $p$ large enough.
+$\frac{|E^{*}|}{p(p-1)^{15}}<\frac{1}{p^{2}}$
+Hence
+$\prod_{p\geqslant N^{9(1-\varepsilon)/200}}\left(1+\frac{E^{*}}{p(p-1)^{15}}\right)\leqslant\prod_{p\geqslant N^{9(1-\varepsilon)/200}}\left(1+\frac{1}{p^{2}}\right)<1+\varepsilon.$
 To sum up,
-$\mathfrak{S}_{1}(n)W(z)<318974.199e^{-\gamma}L^{-1}$
+$\mathfrak{S}_{1}(n)W(z)<16557.733e^{-\gamma}L^{-1}$
 then
-$\sum_{m\sim U\atop(m,P(z))=1}R(m)<bV^{4}L^{-8},$
-where $b=3702014554$.
+$\sum_{m\sim U\atop(m,P(z))=1}R(m)<bU^{4}V^{8}L^{-16},$
+where $b=80947432211.141$.
 Naturally, for $n=0$,
@@
 therefore
-$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)<bV^{4}L^{-8}.$
+$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)<bU^{4}V^{8}L^{-16}.$
 We have proved Lemma 8.
@@
 By the prime number theorem, for
-$U=\left(\frac{N}{32(1+\delta)}\right)^{1/4},\quad V=U^{5/6},$
+$U=\left(\frac{N}{64(1+\delta_{0})}\right)^{1/4},\quad V=U^{7/8},$
 we have
-(4.1) $\begin{split}\sum_{N/9<\ell\leqslant N}r(\ell)&\geqslant\sum_{p_{1}\sim U}1\sum_{p_{2}\sim U}1\sum_{p_{3}\sim V}1\sum_{p_{4}\sim V}1\geqslant(1-\varepsilon)\frac{U^{2}V^{2}}{\log^{2}U\log^{2}V}\\ &\geqslant\left(\frac{96}{5}\right)^{2}(1-\varepsilon)U^{2}V^{2}L^{-4}.\end{split}$
+(4.1) $\begin{split}\sum_{N/9<\ell\leqslant N}r(\ell)&\geqslant\sum_{p_{1}\sim U}1\sum_{p_{2}\sim U}1\sum_{p_{3}\sim U}1\sum_{p_{4}\sim U}1\sum_{p_{5}\sim V}1\sum_{p_{6}\sim V}1\sum_{p_{7}\sim V}\sum_{p_{8}\sim V}\\ &\geqslant(1-\varepsilon)\frac{U^{4}V^{4}}{\log^{4}U\log^{4}V}\geqslant\left(\frac{128}{7}\right)^{4}(1-\varepsilon)U^{4}V^{4}L^{-8}.\end{split}$
 then by Cauchy’s inequality and Lemma 8, we have
-(4.2) $\begin{split}\left\{\sum_{N/9<\ell\leqslant N}r(\ell)\right\}^{2}&\leqslant\left\{\sum_{\begin{subarray}{c}N/9<\ell\leqslant N\\ r(\ell)>0\end{subarray}}1\right\}\left\{\sum_{N/9<\ell\leqslant N}r^{2}(\ell)\right\}\\ &\leqslant bV^{4}L^{-8}\left\{\sum_{N/9<\ell\leqslant N}1\right\}.\end{split}$
+(4.2) $\begin{split}\left\{\sum_{N/9<\ell\leqslant N}r(\ell)\right\}^{2}&\leqslant\left\{\sum_{\begin{subarray}{c}N/9<\ell\leqslant N\\ r(\ell)>0\end{subarray}}1\right\}\left\{\sum_{N/9<\ell\leqslant N}r^{2}(\ell)\right\}\\ &\leqslant bU^{4}V^{8}L^{-16}\left\{\sum_{N/9<\ell\leqslant N}1\right\}.\end{split}$
 From (4.1), (4.2), we have
-$\sum_{N/9<\ell\leqslant N\atop r(\ell)>0}1>\frac{(1-\varepsilon)^{2}}{b}\biggl(\frac{96}{5}\biggr)^{4}U^{4}>\frac{1}{27241.64}N.$
+$\sum_{N/9<\ell\leqslant N\atop r(\ell)>0}1>\frac{(1-\varepsilon)^{2}}{b}\biggl(\frac{128}{7}\biggr)^{8}U^{4}>\frac{1}{414.465}N.$
 The proof of our Theorem is now complete.